HEDGING (10 Marks) Part A (5 Marks)

An oil refiner purchases ·crude oil, which it then refines (i.e. processes) into finished products like gasoline and diesel. Does the refiner have a long or short exposure to crude oil? How about gasoline (i.e. long or short_ exposure)? What hedging positions would the refiner take with respect to crude oil and gasoline if it wants to increase the certainty of its cash flows?

Part B (5 Marks)

CanTruck, a Canadian truck manufacturer, will be delivering a large shipment to a French firm in 1 year. CanTruck expects to receive a payment of 10 million Euros at that time. Currently the spot rate is $1.50 (CON/€) and the 1-year forward rate is $1.60 (CDN/€).

  1. What risks does CanTruck face with the sale of the fleet of trucks?

  2. Describe how CanTruck can hedge the currency risk.

  3. Describe CanTruck's profit or loss on the hedge if-the actual spot rate in 1 year is:

    1. $1.40 per Euro

    2. $1.80 per Euro

  4. Given your answers to (a) and (b), should CanTruck hedge?
Answer #1

Part A: Crude oil is used by refinery as a major Raw Material and he needs to purchase crude oil to process it into finished product, so refiner needs regular supply of crude oil and wants to hedge the price risk that can be done by taking long exposure in crude oil futures as refiner is having upside price risk in crude oi prices. so in this case refiner will be having short exposure in crude oil that can be hedged by taking long position in crude oil futures .

If we talk about gasoline this the finished product for refiner and he has supply gasoline, in this case refiner is facing downside price risk in gasoline as he is having long exposure in gasoline that can be hedged by taking a short position in gasoline to make the cash flow stable.

Part B: CanTruck is exporting trucks to french firm where he is going to receive his payments of 10 Million Euros.

in this case euro is foreign currency and CAD is domestic currency, exporters receives payment in domestic currency (CAD in this case) so depreciation of EURO against CAD is an risk for Canadian Exporter he needs to take short position in Euro/CAD to hedge the depreciation risk against CAD.

Spot Rate: 1.5 EURO/CAD

1 Year forward rate: 1.6 Euro/CAD

payment receivable 10000000 EURO

to hedge the currency exchange risk he needs to short 10000000 EURO @ 1.6 Euro/CAD

He receives 16000000 CAD against 10000000 EURO

After 1 Year

(a) If Exchange rate is 1.4 per Euro (CAD Appreciated against EURO)

From French firm he receives 1000000 Euros and converts @ 1.4 CAD he Receives $14000000

Loss due to depreciation of EURO Against CAD:

Spot at the time of Deal 1.5* 1000000 = $15000000

Spot At the time of Payment 1.4*1000000 = $14000000

Loss= 15000000-14000000 = ($1000000)

From SHORT Forward contract He needs to pay 1000000 Euro which he can buy @1.4 and pays $14000000

Profit from Forward contract 16000000-14000000 = $2000000

Overall P&L From Hedging = $2000000-$100000 = $100000 Profit

(b) If Exchange rate is 1.8 per Euro (CAD Depreciated against EURO)

From French firm he receives 1000000 Euros and converts @ 1.8 CAD he Receives $18000000

Profit due to appreciation of Euro against CAD

Spot at the time of Deal 1.5* 1000000 = $15000000

Spot At the time of Payment 1.8*1000000 = $18000000

Profit = 18000000-14000000 = $3000000

From SHORT Forward contract He needs to pay 1000000 Euro which he can buy @1.8 and pays $18000000

Loss from Forward contract 16000000-18000000 = ($2000000)

Overall P&L From Hedging = $3000000-$200000 = $100000 Profit

Cantruck should hedge this position.